OpenStudy (anonymous):

\[(0.91)^{n−1}=0.998\]

OpenStudy (mr.math):

I suppose you want to find n. By taking ln of both sides we get \[0.91^{n-1}=0.998 \implies (n-1)\ln(0.91)=\ln(0.998)...\] You should be able to take it from here.